### v1

# Occlusion-Aware Seamless Segmentation

Abstract
Panoramic images can broaden the Field of View (FoV), occlusion-aware prediction can deepen the understanding of the scene, and domain adaptation can transfer across viewing domains. In this work, we introduce a novel task, Occlusion-Aware Seamless Segmentation (OASS), which simultaneously tackles all these three challenges. For benchmarking OASS, we establish a new human-annotated dataset for Blending Panoramic Amodal Seamless Segmentation, i.e., BlendPASS. Besides, we propose the first solution UnmaskFormer, aiming at unmasking the narrow FoV, occlusions, and domain gaps all at once. Specifically, UnmaskFormer includes the crucial designs of Unmasking Attention (UA) and Amodal-oriented Mix (AoMix). Our method achieves state-of-the-art performance on the BlendPASS dataset, reaching a remarkable mAPQ of and mIoU of . On public panoramic semantic segmentation datasets, i.e., SynPASS and DensePASS, our method outperforms previous methods and obtains and in mIoU, respectively. The fresh BlendPASS dataset and our source code will be made publicly available at OASS.
Keywords:
Panoramic Scene Understanding Amodal Segmentation1 Introduction
Panoramic imaging has advanced significantly [36, 87], allowing the capture of high-quality 360° images with minimalist optical systems [35] suitable for a wide variety of omnidirectional vision applications [1, 20]. Concurrently, panoramic scene understanding has advanced in many areas, such as dense visual prediction [65, 69, 49, 88], holistic scene understanding [70, 92], and panoramic scene segmentation [78, 58, 81, 30]. On the other side, amodal perception [57], a fundamental aspect of human vision that forms the basis of our understanding and interpretation of the world, motivates occlusion-aware amodal prediction [55, 61, 2] aimed at achieving recognition of an object and its complete spatial extent. These diverse research endeavors converge toward the common goal of achieving more comprehensive visual perception and understanding.
To unify the aforementioned scene understanding in a seamless form and advance a more comprehensive perception, we introduce OASS (Occlusion-Aware Seamless Segmentation). As illustrated in Fig. 1, the OASS task offers threefold benefits while posing three significant challenges: (1) Panoramic images offer a broader Field of View (FoV), unmasking the narrow FoV of the pinhole imagery, e.g., from 95° to 360°, as compared between and of Fig. 1(a). However, panoramas introduce severe distortions compared to pinhole images, which will lead to a significant performance degradation [94, 98]. (2) Amodal prediction [2], in contrast to segmentation limited in visible areas [77, 22], can unmask the occlusion in the space perspective. For example, compared between and of Fig. 1(a), the occluded pedestrian can be completely recognized by using amodal prediction. Predicting the complete mask of occluded objects, i.e., occlusion reasoning, is important to enhance the spatial-wise understanding capacity [9, 90, 41, 37]. (3) Unsupervised Domain Adaptation (UDA) [62, 96] in is capable of addressing the need for expensive training data, unmasking domain gaps from label-rich pinhole to label-scare panoramic imagery.
Previous approaches [94, 98, 5, 61, 55] are proposed to address the above challenges separately, resulting in sub-optimal and non-seamless solutions. For example, as the comparison depicted in Fig. 2, methods [94, 98] proposed to address panoramic semantic segmentation cannot handle the object occlusion, whereas methods [61, 55, 5] proposed for amodal segmentation cannot generalize well to panoramic imagery and the FoV occlusion remains unsolved. To address the challenges seamlessly, we propose a novel unmasking transformer framework called UnmaskFormer. The novelty is three-fold: (1) An Unmasking Attention (UA) constructed by a self-attention and an enhanced pooling layer for occlusion prediction, is proposed to unmask the object occlusions within the whole UnmaskFormer framework. (2) We delve deeper into the design of the Deformable Patch Embedding (DPE) [94] and find an alternative yet better solution for addressing the image distortion of panoramas at different stages within the transformer model. (3) We propose an Amodal-oriented Mix (AoMix) method that aims to seamlessly integrate the pinhole and panoramic domains, addressing the challenges posed by inconsistent cross-domain scenes. This method also enhances the model’s capacity to reconstruct invisible regions of occluded objects, ultimately allowing it to unmask the occlusion of perspective within a scene. Based on these crucial designs, our UnmaskFormer can better handle different panorama-based scene understanding tasks, especially solving object occlusion and image distortion. As shown in Fig. 1(b), our UnmaskFormer manages to attain striking performances across Semantic (mIoU), Instance (mAP), Amodal Instance (mAAP), Panoptic (mPQ), and Amodal Panoptic Segmentation (mAPQ) tasks via one model.
In the literature, there is a lack of datasets or benchmarks suitable for learning and evaluating Occlusion-Aware Seamless Segmentation (OASS) methods. To facilitate this evaluation, we spent a large effort to collect and manually annotate a dataset for Blending Panoramic Amodal Seamless Segmentation (BlendPASS). There are panoramic images with 360° FoV and a resolution, captured in street scenes from cities worldwide, featuring both driving and walking perspectives. BlendPASS facilitates the learning optimization of segmentation adaptation in an unsupervised fashion, which unfolds as efficient for dense prediction in panoramic imagery. To establish the evaluation benchmark for the OASS task, panoramic images have been manually annotated precisely at the pixel level.
Extensive experiments are conducted on our proposed BlendPASS dataset and other panoramic datasets. Our UnmaskFormer achieves state-of-the-art performance on BlendPASS, reaching a remarkable mAPQ of and mIoU of . On the panoramic semantic segmentation datasets SynPASS [95] and DensePASS [53], our method outperforms the previous best method and obtains and in mIoU, respectively. The significant improvements and results prove the effectiveness of the proposed UnmaskFormer framework in addressing the panorama-based scene understanding.
In this work, we propose contributions as follows:
We introduce a new panorama-based segmentation task, i.e., Occlusion-Aware Seamless Segmentation (OASS), aiming at unmasking the narrow field of view, unmasking the occlusion of perspective, and unmasking the gap of domain in a seamless manner.
To address the OASS task, we propose an UnmaskFormer framework with distortion- and occlusion-aware designs in a transformer-based architecture.
An Amodal-oriented Mix (AoMix) method is tailored for improving unsupervised domain adaptation and overcoming the challenges of diverse occlusions in bridging the gap between pinhole and panoramic domains.
A new panoramic dataset BlendPASS is created and manually annotated for benchmarking the blending of panoramic amodal seamless segmentation.
2 Related Work
Domain adaptive panoramic segmentation. Segmentation on wide-FoV fisheye images [13, 66, 85, 63, 86] and panoramic images [23, 43, 80, 83, 100] enables holistic understanding of 360° surroundings [79]. To address the scarcity of annotations, researchers have revisited wide-FoV segmentation from the perspective of unsupervised domain adaptation [38, 32, 84, 82, 67] by leveraging rich training sources from conventional narrow-FoV data. A variety of adaptation methods including self-training [28, 29, 103, 27, 75, 91] and adversarial training [53, 51] methods are studied for panoramic segmentation. In this line, P2PDA [93] employs attention-regulated uncertainty-aware adaptation. Trans4PASS [94, 95] introduces distortion-aware transformers and mutual prototypical adaptation with SAM [40]. DPPASS [99] presents a dual-path solution to overcome style and format gaps. DATR [98] captures neighboring distributions without any geometric constraints. Moreover, panoramic panoptic segmentation [18, 34, 54] is investigated by using techniques like contrastive learning for harvesting generalization benefits [33]. In this work, we aim to enable occlusion-informed understanding with both FoV-wise and spatially amodal reasoning of urban scenes.
Amodal scene segmentation. Amodal instance segmentation is a derivative task of instance segmentation [25, 12, 4, 68, 71, 46, 16], to predict visible regions of objects and their occluded regions simultaneously. Li et al. [42] introduce the concept of amodal instance segmentation, achieved through iterative regression. Zhu et al. [102] derive two amodal segmentation datasets from BSDS [3] and COCO [48]. Qi et al. [61] derive the amodal version of KITTI [21], introducing an independent occlusion classification branch. SLN [97] incorporates a semantics-aware distance map to implement amodal segmentation. ORCNN [17] calculates visible masks to infer occlusion masks. Further, shape and contour priors are frequently leveraged for refining amodal segmentation [19, 6, 45, 76, 44]. Aside from amodal instance segmentation, amodal semantic understanding [102, 31, 5] is an important direction. Along this line, [55, 56] propose a fusion of semantic and amodal instance segmentation. Sekkat et al. [64] uses the CARLA simulator [14] to create a virtual multi-task amodal perception dataset. Different from these works, we intend to achieve occlusion-aware seamless segmentation, which breaks the occlusion limits in terms of both field-of-view and in-field object occlusions.
3 Established Benchmark
3.1 Overview of the Benchmark
In this work, we establish a novel benchmark specifically designed for Occlusion-Aware Seamless Segmentation (OASS). In particular, we address OASS from the perspective of UDA by adapting from the label-rich pinhole domain to the label-scarce panoramic domain. 360° panoramas have a wide FoV and many small objects in the images, which exaggerate the costs of creating pixel-wise annotations in unconstrained surroundings. Our benchmark overcomes the scarcity of amodal segmentation testbeds in panoramic imagery, while addressing three aforementioned challenges: how to address the severe distortions of panoramas when unmasking the narrow field of view, how to predict the full segmentation of objects when unmasking the occlusion of perspective, and how to facilitate optimization of segmentation adaptation when unmasking the gap of domain.
Our objective with this benchmark is to provide a comprehensive evaluation of methods capable of performing both FoV-wise and spatially occlusion-aware seamless segmentation. We extend the metrics proposed by [39] to the full regions incorporating pixels of the occluded objects. The benchmark metrics cover three aspects: Intersection over Union (IoU) for semantic segmentation, Average Precision (AP) and Panoptic Quality (PQ) for instance and panoptic segmentation, Amodal Average Precision (AAP) and Amodal Panoptic Quality (APQ) for amodal instance and amodal panoptic segmentation.
| Person | Rider | Car | Truck | Bus | Train | Motorcycle | Bicycle | Total | ||
189 | 6 | 909 | 42 | 18 | 1 | 83 | 38 | 1286 | |
613 | 12 | 842 | 38 | 24 | 2 | 71 | 72 | 1674 | |
| Total | 802 | 18 | 1751 | 80 | 42 | 3 | 154 | 110 | 2960 |
3.2 BlendPASS
We introduce a novel dataset for Blending Panoramic Amodal Seamless Segmentation (BlendPASS) tailored for OASS. BlendPASS comprises an unlabeled training set of panoramic images for optimizing domain adaptation and a labeled test set of panoramic images. These images are captured from panoramic cameras in driving scenes across cities for the training set and cities for the test set, all at a resolution of pixels. As depicted in Fig. 3(a), we have provided pixel-wise annotations for five distinct visual segmentation tasks, which greatly extends the semantic labels from DensePASS [53]. These annotations cover categories that align with the Cityscapes [11] and are further categorized into Stuff (road, sidewalk, building, wall, fence, pole, light, sign, vegetation, terrain, and sky) and Thing (person, rider, car, truck, bus, train, motorcycle, and bicycle). To ensure the precision of annotations and the rational handling of occluded object parts, we meticulously conduct manual annotation for the test set, with three annotators following the “independent annotating cross-checking voting” workflow, as outlined in [55]. Finally, as illustrated in Fig. 3(b), objects are annotated in the Thing class, with of these objects exhibiting occlusion.
3.3 KITTI360-APSBlendPASS
For the labeled source in our UDA benchmark for OASS, we employ the available KITTI360-APS dataset [55] designed for amodal panoptic segmentation. This dataset is an extension of KITTI360 [47] and includes annotations for inmodal/amodal instance and panoptic segmentation. The images in KITTI360-APS are captured using pinhole cameras from cities, with a resolution of . After our careful examination, a total of annotated images are accessed. These annotations encompass Stuff (road, sidewalk, building, wall, fence, pole, traffic sign, vegetation, terrain, and sky) classes and Thing (car, pedestrians, cyclists, two-wheeler, van, truck, and other vehicles) classes. In our benchmark, we further process the annotations from BlendPASS to ensure class alignment with KITTI360-APS.
For KITTI360-APSBlendPASS under OASS, besides the challenges mentioned in Sec. 3.1, as illustrated in Fig. 4, the source-target domains exhibit significant differences in terms of the number of objects per image and the class distribution. These differences present a greater challenge to UDA models.
4 Methodology
In this section, we describe the proposed method UnmaskFormer tailored for OASS, as illustrated in Fig. 5. We start by introducing the principal architecture of UnmaskFormer, featuring Unmasking Attention (UA) blocks crafted for distortion- and occlusion-aware reasoning in OASS. Subsequently, we elaborate on our cross-domain adaptation approach, leveraging the Amodal-oriented Mix (AoMix) method to enhance occlusion-aware knowledge transfer.
4.1 UnmaskFormer
Architecture. For the OASS task, the model simultaneously addresses objectives including semantic, instance, amodal instance, panoptic, and amodal panoptic segmentation. Similarly to panoptic segmentation, as illustrated in Fig. 5, we decompose the base model into a feature extractor and three branches, i.e., semantic, instance, and amodal instance, to accomplish OASS goals. The semantic branch predicts the per-pixel semantic category of the input image, while the instance and amodal instance branches output class-agnostic object localization predictions. It is noteworthy that, while the instance branch has both top-down [62] and bottom-up [96] decoders, the contour-based bottom-up decoder is impractical for the amodal instance branch in the context of OASS. This limitation arises from situations where a single pixel is associated with multiple objects simultaneously. In contrast, the proposals-based top-down decoder typically follows Mask R-CNN [26], employing Region Proposal Networks (RPN) to predict candidate objects. This methodology facilitates simultaneous and interference-free segmentation of multiple objects. Hence, in UnmaskFormer, we adopt the top-down decoder for the instance branch and the occlusion-aware amodal instance branch.
For the final output of the UnmaskFormer, we construct an Occlusion-Aware Fusion (OAFusion) module to process outputs from three branches, generating five segmentation maps at once for OASS. The semantic segmentation is directly predicted by the semantic branch. In instance and amodal instance segmentation, the semantic label of class-agnostic instance or amodal instance mask is determined by the majority-voting rule of the predictions from the semantic branch. For amodal segmentation, only regions where the current object does not overlap with other objects are considered. The panoptic and amodal panoptic segmentation are generated by fusing the semantic segmentation with instance and amodal instance segmentation, respectively.
UA-based backbone. One of challenges in OASS is to address both image distortion and object occlusion at the same time. For this end, we construct a UA-based backbone by rearranging previous Deformable Patch Embedding (DPE) and adding an effective pooling layer, as shown in Fig. 5. Specifically, for an input target image or features between different stages, the original DPE calculates the adaptive offsets and performs the patchifying process. Contrary to inserting DPE in the early stage [94], we perform a new interleaving arrangement of DPE layers in Stages 2 and 4, which can provide reinforced distortion-aware modeling for the whole UnmaskFormer framework. More analysis will be presented in Sec. 5.4.
Apart from the distortion-aware design, we design the UA block by combining the self-attention layer and the enhanced pooling layer. As shown in Fig. 5, after the self-attention layer, the feature is forwarded to calculate the pooling query , the key and the value . After having the self-attended pooling feature , a sigmoid function is applied to calculate the occlusion-aware mask . Then, the pool-attended occlusion-aware mask is used to perform a dot product operation with feature to obtain the occlusion-aware feature . After the UA block, the occlusion-aware feature is further forwarded to an MLP layer as self-attention blocks [77, 94].
4.2 Cross-Domain Adaptation
UDA strategy. In the OASS setting, the model is assigned to adapt from a labeled source domain to an unlabeled target domain for multiple segmentation tasks, which are images from pinhole and panoramic cameras respectively. To address the domain gap arising from dissimilar data distributions between the source pinhole and target panoramic domains, a comprehensive loss is utilized for training the overall network :
where and represent the source domain supervision loss and the target domain adaptation loss, respectively. is utilized for supervision to learn fundamental segmentation capacity through labeled source samples , whereas the leverages unlabeled target images to enhance segmentation in the target panoramic scenario. includes semantic loss, instance loss, and amodal instance loss, each originating from respective branches. The semantic branch employs cross-entropy loss to assign each pixel to its respective category. As for the instance and amodal instance branches, we follow the conventional approach from Mask R-CNN [26], employing supervision using proposals-based bounding boxes and instance-level masks.
To tackle the challenge posed by the absence of labels in the target domain, we employ a self-training strategy to facilitate the model’s adaptation from the source to the target domain. Self-training methods [28, 29, 27] typically utilize the target predictions as pseudo-labels for training. However, due to the noisy predictions in the early stage, resulting in low-confidence pseudo-labels, we adopt a confidence estimation mechanism that assigns weights of pseudo-labels-based self-training loss . Additionally, we incorporate the Mean-Teacher framework widely adopted in UDA [27, 72, 62] to further enhance the quality of the pseudo-labels. Thus, can be expressed as follows:
where
The parameters of the teacher model are updated using the parameters of the student model by the exponential moving average at each iteration :
Amodal-oriented Mix (AoMix). To address the challenges of diverse occlusion cases and inconsistent cross-domain scenes, we propose a new method AoMix, which processes images from two domains alongside self-training. Initially, we reconstruct the masked source images based on a random source amodal annotations . Subsequently, we follow the widely-used class-mix strategy [72, 27] in UDA segmentation tasks to generate masked mixed images , incorporating information from both the source and target domains.
Specifically, we randomly sample an amodal instance mask sequence of an image from the current batch, and subject it to random scaling and peripheral random padding to generate a new amodal instance mask sequence with the object of various positions and sizes. Then, a random binary mask containing multiple objects is produced by summing these masks . This operation can be expressed as follows:
where denoted as a step function. For a source image , we sum the corresponding amodal instance masks to obtain a binary mask of Thing region. Using the random binary mask and the Thing region binary mask , we fill the source image to reconstruct the masked source image .
Amodal-oriented masked image modeling aims to enhance the model’s ability to reconstruct object regions obscured by realistic object shapes, enabling it to learn information about invisible parts in the scene. To adapt the model to unlabeled target scenes, we randomly sample half of the semantic classes from a source image and paste the corresponding pixels of the masked source image onto a target image based on the source semantic map, creating a new masked mixed image . This approach effectively transfers the model’s capability tailored for distortion- and occlusion-aware to the target panoramic domain.
5 Experiments
5.1 Experiment Setups
Following DAFomer [27], we train UnmaskFomer using the AdamW optimizer [50]. The learning rate is set to , with a weight decay of and linear warmup for iterations followed by polynomial decay. The model is trained on a batch size of with a crop size of for iterations. The EMA decay is set as . We use the same data augmentation parameters as DACS [72] and set the RCS temperature to to enhance sample frequency for rare classes of the source domain. Moreover, we adopt the ImageNet feature distance loss from DAFormer with a weight of . For instance and amodal instance branches, we set all loss weights to . Our experiments are conducted on an NVIDIA Tesla V100 GPU and implemented using PyTorch.
| Task | UDA Method | road | sidewalk | building | wall | fence | pole | traffic-light | traffic-sign | vegetation | terrain | sky | pedestrians | cyclists | car | truck | other-vehicles | van | two-wheeler | Metric |
| SS | DATR [98] | 71.87 | 27.24 | 70.59 | 22.76 | 35.98 | 23.91 | 00.00 | 04.52 | 77.06 | 37.14 | 80.06 | 51.20 | 02.24 | 70.15 | 08.94 | 06.00 | 11.02 | 27.79 | 34.91 |
| Trans4PASS [94] | 72.79 | 33.80 | 78.38 | 33.53 | 37.07 | 26.56 | 05.31 | 05.35 | 77.43 | 37.91 | 84.73 | 57.48 | 04.51 | 76.58 | 19.04 | 17.03 | 12.66 | 51.70 | 40.66 | |
| UniDAPS [96] | 72.27 | 29.20 | 75.78 | 33.91 | 38.92 | 25.94 | 11.37 | 07.51 | 77.61 | 37.40 | 81.40 | 47.56 | 02.95 | 75.59 | 21.29 | 03.39 | 01.31 | 48.85 | 38.46 | |
| EDAPS [62] | 74.41 | 35.91 | 76.98 | 36.45 | 40.35 | 28.02 | 16.08 | 05.13 | 78.10 | 39.50 | 82.28 | 55.45 | 03.25 | 74.38 | 06.72 | 14.09 | 05.18 | 50.87 | 40.17 | |
| Source-Only | 72.97 | 29.14 | 82.04 | 31.19 | 31.37 | 17.98 | 00.00 | 15.97 | 74.13 | 33.74 | 88.66 | 50.24 | 04.02 | 80.42 | 18.02 | 11.04 | 04.57 | 50.17 | 38.65 | |
| UnmaskFormer | 76.54 | 37.82 | 77.06 | 34.71 | 44.05 | 28.27 | 17.80 | 02.76 | 78.70 | 41.68 | 84.98 | 57.34 | 06.01 | 80.60 | 23.47 | 21.67 | 18.80 | 53.56 | 43.66 | |
| APS | DATR [98] | 51.82 | 09.15 | 59.90 | 11.93 | 11.98 | 01.97 | 00.00 | 03.91 | 64.61 | 14.05 | 70.40 | 11.09 | 00.00 | 39.30 | 00.00 | 03.16 | 10.07 | 01.30 | 20.26 |
| Trans4PASS [94] | 53.91 | 14.12 | 69.39 | 19.15 | 11.77 | 03.77 | 00.00 | 05.15 | 67.63 | 16.02 | 77.41 | 15.30 | 04.24 | 41.06 | 06.58 | 00.00 | 00.00 | 07.35 | 22.94 | |
| EDAPS [62] | 54.88 | 17.04 | 66.86 | 18.75 | 14.47 | 05.75 | 04.04 | 04.64 | 68.20 | 16.04 | 72.76 | 19.01 | 00.00 | 36.73 | 05.77 | 04.38 | 00.00 | 07.20 | 23.14 | |
| Source-Only | 57.11 | 14.21 | 73.58 | 15.49 | 07.59 | 00.67 | 00.00 | 10.40 | 58.30 | 12.39 | 83.14 | 15.23 | 00.00 | 40.31 | 03.83 | 00.00 | 00.00 | 06.08 | 22.13 | |
| UnmaskFormer | 61.84 | 24.72 | 66.81 | 20.77 | 15.80 | 05.25 | 04.29 | 03.26 | 69.02 | 18.35 | 79.44 | 20.48 | 03.14 | 44.56 | 12.81 | 11.25 | 00.00 | 16.64 | 26.58 |
5.2 Results of OASS on KITTI360BlendPASS
We deliver the OASS results on the KITTI360-APSBlendPASS benchmark in Tab. 1. Representative UDA panoptic segmentation methods UniDAPS [96] and EDAPS [62] with our amodal extensions are benchmarked. Further, state-of-the-art panoramic semantic segmentation models Trans4PASS [94] and DATR [98] with distortion-adaptive capacities are compared. It is worth noting that UniDAPS does not support occlusion-aware reasoning with a contour-based decoder.
As shown in Tab. 1, compared with the best-performing UDA panoptic segmentation model EDAPS [62], UnmaskFormer outstrips it by clear margins of in mIoU and in mAPQ. Compared to the panoramic segmentation model Trans4PASS [94], our method also yields significant gains. Delving into the OASS results in semantic- and amodal panoptic segmentation, all benchmarked methods suffer from accurately predicting the full segmentation in the cases considering occlusions, in particular on small objects, evidenced by the unsatisfactory scores in class-wise APQ. Yet, UnmaskFormer harvests great improvements in contrast to the Source-Only model, e.g., on safety-critical pedestrians, truck, and two-wheeler. Finally, under the challenging occlusion scenarios in unstructured surroundings, UnmaskFormer clearly stands out and leads to state-of-the-art OASS scores of in mIoU and in mAPQ. In Fig. 6, we visualize OASS results produced by our UnmaskFormer and state-of-the-art methods DATR [98], Trans4PASS [94], EDAPS [62]. Compared to them, UnmaskFormer excels in segmenting occluded objects and realistically reconstructing vehicle shapes thanks to our UA-based backbone and AoMix design.
| Stuff: | road | sidew. | build. | wall | fence | pole | tr.light | tr.sign | veget. | terrain | sky |
| Thing: | pedes. | cyclists | car | truck | ot.veh. | van | tw.whe. |
| UDA Method | A | pedestrians | cyclists | car | truck | other-vehicles | van | two-wheeler | Metric |
| DATR [98] | 14.21 | 00.00 | 31.15 | 07.55 | 03.73 | 00.39 | 03.57 | 08.66 | |
| ✓ | 13.11 | 00.03 | 30.60 | 06.87 | 04.73 | 01.67 | 03.78 | 08.68 | |
| Trans4PASS [94] | 16.52 | 00.03 | 32.23 | 10.19 | 05.31 | 00.16 | 05.62 | 10.01 | |
| ✓ | 15.95 | 00.21 | 31.71 | 08.34 | 06.01 | 00.35 | 06.41 | 09.85 | |
| UniDAPS [96] | 02.31 | 00.00 | 11.25 | 06.17 | 02.80 | 00.00 | 01.50 | 03.43 | |
| ✓ | n.a. | n.a. | n.a. | n.a. | n.a. | n.a. | n.a. | n.a. | |
| EDAPS [62] | 16.61 | 00.04 | 30.83 | 06.49 | 11.42 | 00.37 | 06.19 | 10.28 | |
| ✓ | 15.78 | 00.09 | 30.02 | 09.00 | 12.20 | 00.36 | 07.35 | 10.68 | |
| Source-Only | 15.79 | 00.02 | 33.88 | 13.54 | 03.58 | 00.19 | 06.78 | 10.54 | |
| ✓ | 15.18 | 00.01 | 33.37 | 12.60 | 03.28 | 00.35 | 06.77 | 10.22 | |
| UnmaskFormer | 17.55 | 00.00 | 35.18 | 14.18 | 02.92 | 00.77 | 07.14 | 11.10 | |
| (Ours) | ✓ | 16.10 | 00.07 | 34.07 | 12.29 | 02.19 | 00.61 | 08.15 | 10.50 |
Tab. 2 additionally presents the results in amodal instance segmentation of our comparative experiments on the benchmark of KITTI360-APSBlendPASS transfer. The instance segmentation performance measured in both AP for visible regions and AAP for full regions are listed. UnmaskFormer reaches a new record of in mAP and reaches an on-par mAAP score of as EDAPS [62] specifically designed for UDA panoptic segmentation. Compared to EDAPS, our UnmaskFormer has large gains on challenging classes like trucks and cars frequently under large occlusions in unconstrained scenes.
5.3 Results of Panoramic Semantic Segmentation
SynPASS dataset. We further investigate the generalization capacity of the proposed UnmaskFormer on panoramic semantic segmentation. In Tab. 4, we conduct a comparison between known convolutional and transformer models for panoramic semantic segmentation on the SynPASS benchmark [95]. Here, the models are learned on the training set of images and evaluated on the validation set of images with diverse weather and illumination conditions. It can be seen that UnmaskFormer equipped with the unmasking attention and the semantic head, produces state-of-the-art performance measured in mIoU, which reaches . Moreover, in each adverse condition, UnmaskFormer yields robust segmentation of the full panoramas on SynPASS.
| Method | Cloudy | Foggy | Rainy | Sunny | Day | Night | ALL |
| Fast-SCNN [59] | 30.84 | 22.68 | 26.16 | 27.19 | 29.68 | 24.75 | 26.31 |
| DeepLabv3+ (MNv2) [7] | 38.94 | 35.19 | 35.43 | 37.73 | 36.01 | 30.55 | 36.72 |
| HRNet (W18Small) [73] | 42.92 | 37.94 | 37.37 | 41.45 | 39.19 | 32.22 | 39.80 |
| PVT (Small) [74] | 40.75 | 36.14 | 34.29 | 40.14 | 37.92 | 28.80 | 37.47 |
| SegFormer (B2) [77] | 46.07 | 40.99 | 40.10 | 44.35 | 44.08 | 33.99 | 42.49 |
| Trans4PASS (Small) [94] | 46.74 | 43.49 | 43.39 | 45.94 | 45.52 | 37.03 | 44.80 |
| UnmaskFormer (Ours) | 48.00 | 43.43 | 43.48 | 47.06 | 45.87 | 36.43 | 45.34 |
DensePASS dataset. As shown in Tab. 4, we further compare UnmaskFormer against efficient panoramic segmentation transformers including DPPASS [99], DATR [98], and Trans4PASS [95], where the models are trained on Cityscapes [11] and tested on the original panoramic semantic segmentation benchmark DensePASS [53]. Compared to the state-of-the-art DATR [98] constituted by distortion-aware attention with parameters, our UnmaskFormer with fewer parameters () attains higher segmentation performance. Our UnmaskFormer achieves in mIoU, which outperforms its counterparts and maintains as a lightweight segmenter. Therefore, UnmaskFormer is suitable for efficiency-relevant panoramic driving scene understanding.
5.4 Ablation Study
To better understand the components of the UnmaskFormer, we conduct ablation studies for the Unmasking Attention (UA)-based backbone, Amodal-oriented Mix (AoMix) method, and Occlusion-aware Fusion (OAFusion) module.
Why UA unmasks occlusions? In Tab. 7, we ablate different patch embedding methods [77, 94] and pooling methods [60]. Compared to the original Patch Embedding (PE) design and the early-stage DPE, we construct the backbone by using a new interleaving arrangement which improves the distortion-aware modeling, yielding the best mAPQ score with a and a gain respectively. Compared to SimPool [60], our occlusion-aware pooling attention better handles object occlusion and obtains higher results in OASS. An example of feature visualization in Fig. 8 shows that our UA can enhance distortion-tolerant predictions of regions where occlusion occurs, highlighted in red color. These results show the effectiveness of the UA backbone in addressing image distortion and object occlusion at the same time.
How AoMix boosts amodal prediction? As shown in Tab. 7, we investigate various approaches to masked image modeling. Compared to using masked images only for the source image (T for S) or the mixed image (T for M), our strategy (AoMix), which utilizes both, achieves the best results. Masking the Thing class regions (AoMix), as opposed to the entire image (W for S&M) or masked by patches (P for S&M), results in improvements with a gain of and in mAPQ respectively. With the carefully devised AoMix, the UnmaskFormer earns significant gains by greatly boosting surrounding parsing seamlessly.
What OAFusion fuse? As illustrated in Fig. 8, a substantial portion of the pedestrian area (in red) is occluded by the car (in blue). Despite the amodal instance branch being capable of predicting the full region of the pedestrian, different fusion approaches yield distinct category classifications. The conventional fusion strategy [62, 10] tailored for panoptic segmentation, given the substantial occlusion of the pedestrian region by the car, misclassifies it as car under the majority voting rule. Conversely, our designed OAFusion ensures accurate classification of the full region as the pedestrian by disregarding overlapping regions.
Finally, as shown in Tab. 7, all components of UnmaskFormer cooperate with each other to achieve the best performance.
5.5 Analysis of Hyperparameters
Analysis of deformable designs. In Fig. 10, we conduct experiments to analyze designs of PE [77] and DPE [94] in our UnmaskFormer model. We obtain three insights: (1) The PE cannot effectively handle the image distortion in OASS, which is in line with the finding in [94]. (2) Using more DPE blocks inside a four-stage model does not ensure optimal performance. (3) Compared with shallow stages, DPE can bring more improvements when acting on deep stages, e.g., using DPE in Stages 2 and 4 (DPE - 2, 4) performs better than in Stages 1 and 3 (DPE - 1, 3). Therefore, we adopt the deformable design of DPE in Stages 2 and 4 to construct the UnmaskFormer as default, which provides the optimal architecture for addressing image distortion and object deformation in OASS.
Analysis of AoMix parameters. In Fig. 10, we conduct experiments to analyze the impact of different scaling parameters in random scaling operation on the model. Excessively large scaling sizes lead to complete occlusion of objects in images, preventing the model from learning original object information. Overly small scaling sizes result in minimal object occlusion, limiting the model’s capacity to make reasonable predictions for occluded regions. Optimal performance is achieved by setting the random scaling range in to , effectively generating diverse masked source images and masked mixed images. This rich variety of amodal-oriented samples bridges the gap between the pinhole and panoramic domains, enhancing the model’s ability to reconstruct invisible regions of occluded objects and boost seamless segmentation.
6 Conclusion
In this work, we have introduced the task of Occlusion-Aware Seamless Segmentation (OASS) for holistic scene understanding. To address OASS, we put forward UnmaskFormer for unmasking the narrow field of view, unmasking the occlusion of perspective, and unmasking the gap of domain seamlessly. We establish the BlendPASS dataset for facilitating the optimization and evaluation of OASS models, as well as fostering future research in the panoramic and panoptic vision field. Experiments on the fresh BlendPASS, as well as public SynPASS and DensePASS benchmarks, demonstrate the effectiveness of the proposed methods.
A Datasets
Dense annotation of BlendPASS. To enhance the accuracy of dense annotations for objects of the Thing class, particularly in representing invisible regions of occluded objects realistically, our labeling process follows the “independent annotation cross-verification voting” workflow. All images are densely annotated by three skilled annotators using the EISeg tool [24] for initial segmentation. Specifically, three annotators densely annotate the full region of each object, and occluded objects are independently annotated by three annotators. Subsequently, cross-verification is conducted among the annotators. In cases where there are slight discrepancies in the annotations of occluded objects, the final annotation is determined through majority voting. For annotations that do not reach a consensus, more iterative annotation processes are employed until an agreement is reached. This annotation workflow aims to ensure the accuracy of dense annotations while maximizing the fidelity of annotations for occluded regions to the true object shape. Annotating panoramas with severe distortion is challenging and extremely time-consuming, requiring approximately 210 minutes per person per image.
SynPASS. The SynPASS dataset [95] is a panoramic semantic segmentation dataset captured via the Carla simulator [14]. It has four weather conditions including sunny, cloudy, foggy, and rainy scenes, together with daytime and nighttime situations. Overall, it has panoramic images with a resolution of corresponding to a full Field of View (FoV) of , divided into training/validation/testing sets of images, respectively.
CityscapesDensePASS. The CityscapesDensePASS benchmark [53] measures the performance of pixel-wise semantic segmentation models learned on labeled pinhole images of Cityscapes and transferred to DensePASS, which has panoramic images for evaluation. The pinhole images have a resolution of and the panoramic images have a resolution of .
KITTI360-APSBlendPASS. We carefully reviewed the accessible annotations for KITTI360-APS [55], and images containing objects of thing classes are available. This dataset serves as the source pinhole domain of our proposed benchmark. For the introduced fresh BlendPASS dataset, considering the scalability of future work, we annotated images containing objects of thing classes in the evaluation set, following the format of Cityscapes [11]. This serves as the validation set for the target panoramic domain of the benchmark. However, due to inconsistent classes between the two datasets, we conducted additional manual processing. Specifically, in the KITTI360-APS, there are valid Stuff and valid Thing classes (while the work [55] claims classes for Stuff, our manual verification confirmed that the traffic light class is an additional usable annotation class). These Stuff classes align with BlendPASS. As for Thing classes, we adjusted the annotations of BlendPASS to align with KITTI360-APS, following the corresponding scheme in Tab. S1. This rough alignment further magnifies the challenges of cross-domain in the benchmark. Finally, for the OASS benchmark, there are a total of aligned Stuff and aligned Thing classes in both domains.
| Dataset | Categories | |||||||
| BlendPASS | Person | Rider | Car | Truck | Bus | Train | Motorcycle | Bicycle |
| KITTI360-APS [55] | Pedestrains | Cyclists | Car | Truck | Other-vehicles | Van | Two-wheeler | Two-wheeler |
B Evaluation Metrics
In the context of the Occlusion-Aware Seamless Segmentation (OASS) benchmark, we employ five metrics, namely Intersection over Union (IoU), Average Precision (AP), Amodal Average Precision (AAP), Panoptic Quality (PQ), and Amodal Panoptic Quality (APQ), to evaluate the model’s performance. We provide a detailed explanation:
IoU. IoU measures the overlap between predicted segment and ground truth segment and is calculated as:
AP. We follow Pascal VOC [15] and COCO [48] and use average precision, which is computed by averaging the ten equally spaced IoU thresholds from to .
AAP. The AAP is an extended metric of AP for Amodal Instance Segmentation, where the ground truth is replaced with amodal segments.
PQ. We adapt the standard panoptic quality metric proposed by [39] and compute it as
where is a set of True Positive matches.
APQ. The APQ is an extended metric of PQ for Amodal Panoptic Segmentation, where the ground truth segments are replaced with amodal segments.
C More Results
C.1 More OASS Results
| Task | UDA Method | road | sidewalk | building | wall | fence | pole | traffic-light | traffic-sign | vegetation | terrain | sky | pedestrians | cyclists | car | truck | other-vehicles | van | two-wheeler | Metric |
| PS | DATR [98] | 50.44 | 09.14 | 59.92 | 11.93 | 11.98 | 01.95 | 00.00 | 03.91 | 64.60 | 14.05 | 70.45 | 12.15 | 00.00 | 38.09 | 00.00 | 03.38 | 00.00 | 01.29 | 19.63 |
| Trans4PASS [94] | 53.93 | 14.12 | 69.39 | 19.16 | 11.77 | 03.77 | 00.00 | 05.15 | 67.62 | 16.02 | 77.41 | 14.60 | 04.09 | 38.23 | 06.91 | 00.00 | 00.00 | 07.19 | 22.80 | |
| UniDAPS [96] | 65.95 | 9.48 | 66.31 | 17.39 | 14.28 | 04.77 | 00.00 | 06.14 | 67.19 | 16.10 | 72.68 | 08.27 | 00.00 | 27.25 | 14.82 | 09.23 | 00.00 | 08.86 | 22.71 | |
| EDAPS [62] | 55.01 | 17.05 | 66.84 | 18.72 | 14.49 | 05.76 | 04.04 | 04.68 | 68.21 | 16.04 | 72.76 | 19.56 | 00.00 | 37.83 | 01.82 | 04.38 | 00.00 | 07.89 | 23.06 | |
| Source-Only | 57.84 | 14.21 | 73.83 | 15.49 | 07.59 | 00.67 | 00.00 | 10.40 | 58.30 | 12.39 | 83.15 | 14.85 | 00.00 | 39.06 | 05.96 | 00.00 | 00.00 | 07.68 | 22.30 | |
| UnmaskFormer (Ours) | 61.73 | 24.72 | 66.80 | 20.75 | 15.81 | 05.22 | 04.29 | 03.26 | 69.02 | 18.35 | 79.44 | 20.90 | 03.45 | 42.96 | 11.26 | 07.64 | 00.00 | 15.98 | 26.20 |
In this work, we include a comprehensive set of sub-tasks, including semantic segmentation, instance segmentation, amodal instance segmentation, panoptic segmentation, and amodal panoptic segmentation for OASS. To provide a thorough analysis of these diverse tasks, we present the per-class accuracy results. Tab. S2 further details the per-class accuracy specifically in panoptic segmentation (PS). To benchmark our model, we conduct a comparative study against previous state-of-the-art methods, namely DATR [98], Trans4PASS [94], UniDAPS [96], and EDAPS [62]. As shown in the experimental results, our model demonstrates a notable Mean-PQ of , surpassing the performance of the previous best model by a significant margin of . A detailed breakdown of per-class performance reveals that our model excels across numerous categories, such as sidewalk, wall, pedestrians, car, and two-wheeler. It is worth noting that the van category, constituting a minor proportion in both source and target datasets, poses a challenge for all models to handle effectively. The superior mPQ achieved by our model proves its effectiveness in addressing the challenges in the panoptic segmentation part of the OASS task.
C.2 More Analysis of Ablation Study
Preliminary of Deformable Patch Embedding. The backbone utilizes a transformer-based structure with a novel arrangement of Deformable Patch Embedding (DPE) [94] layers. DPE aims to capture local geometric variations caused by image distortion. Given an input image or features (where and represent the resolution and the number of channels), DPE calculates adaptive offsets for each patch:
where denotes the patch index, is the offset prediction function, is a hyperparameter controlling the maximum allowed offset.
Analysis of UA-based Backbone. To improve the modeling capacity of our OASS model, we take not only the deformable patchifying but also the unmasking into account. Specifically, the technical designs involve multiple perspectives. 1) Interleaved DPE arrangement: Unlike prior work [94] that inserts DPE only in the initial stage, this approach proposes a novel interleaving arrangement. DPE layers are strategically placed within Stages 2 and 4 of the backbone architecture. This design reinforces the model’s ability to capture distortion throughout the processing pipeline. 2) Unmasking Attention (UA) for occlusion handling: In addition to addressing distortion, the method incorporates an Unmasking Attention (UA) block that is enhanced by adding a simple yet effective pooling layer. 3) Combining self-attention and enhanced pooling: The UA block leverages both a self-attention layer and an improved pooling mechanism. Having the self-attended pooling feature , a sigmoid function is applied to calculate the occlusion-aware mask . This mask highlights regions likely affected by occlusion. 4) Incorporating occlusion awareness: The mask is subsequently used to perform element-wise multiplication with the original feature map, resulting in an occlusion-aware feature. The occlusion-aware feature is further processed by an MLP layer. Based on these crucial designs, our UA-based backbone can address both the image distortion and object occlusion.
Analysis of AoMix. To provide additional amodal-oriented source priors and mixed image samples to enhance the adaptation, we utilize amodal instance masks to mask input images within the AoMix module. An example of the amodal-oriented masked image modeling process for the source image is illustrated in Fig. S1. This method aims to enhance the model’s ability to reconstruct object regions obscured by realistic object shapes, enabling it to learn information about invisible parts in the scene. As shown in Tab. 6, we conducted an ablation study on the manner and strategy of AoMix. 1) T for S, T for M vs AoMix: Using the amodal-oriented masked image in both the source and mixed images enables the model to accurately segment the full regions of occluded objects in the source domain and facilitates better model adaptation to the target panoramic domain. 2) P for SM vs AoMix: It is noteworthy that we tried to use random patches instead of amodal instance masks to mask images. Although the mIoU score remained largely unchanged, the mAPQ score witnessed a significant drop of 3.3%, affirming that masks with real shapes provide better guidance for learning occlusion-ignored segmentation ability. 3) W for SM vs AoMix: Compared with applying amodal masks to all regions of the image, our method only applies to the Thing class region, resulting in superior performance. This is because our method aligns more closely with the real-world scenario where object occlusion of the Thing class is often raised by other objects of the Thing class.
D More Visualization Results
D.1 More Visualization Results of OASS
We showcase visualization results for semantic segmentation and panoptic segmentation in Fig. S2 and Fig. S3. These examples show that UnmaskFormer achieves outstanding performance on other tasks within the OASS benchmark. In addition to addressing the occlusion of perspective, UnmaskFormer can unmask the narrow field of view and the gap of domain. As depicted in Fig. S2 for panoptic segmentation, UnmaskFormer surpasses other methods [98, 94, 96, 62] and successfully detects more pedestrians. Compared to the contour-based UniDAPS [96], UnmaskFormer segments the Thing objects with more complete and rational shapes. For the category determination of instance masks, existing methods typically rely on fusing results from the semantic branch. Moreover, the visualization results for semantic segmentation in Fig. S3 show that UnmaskFormer achieves more accurate semantic classification results and demonstrates robustness to distortions introduced by wide-FoV panoramic images.
| Stuff: | road | sidew. | build. | wall | fence | pole | tr.light | tr.sign | veget. | terrain | sky |
| Thing: | pedes. | cyclists | car | truck | ot.veh. | van | tw.whe. |
| Stuff: | road | sidew. | build. | wall | fence | pole | tr.light | tr.sign | veget. | terrain | sky |
| Thing: | pedes. | cyclists | car | truck | ot.veh. | van | tw.whe. |
D.2 More Visualization Results on SynPASS
As shown in Fig. S4, we conduct qualitative analyses of panoramic semantic segmentation on the SynPASS dataset. The objective of this analysis is to evaluate the robustness of various methods under diverse weather conditions. Specifically, we examine four distinct weather types, namely cloudy, foggy, rainy, and sunny, to assess the adaptability of the models. From top to bottom in Fig. S4, they are the input image, the segmentation results of Trans4PASS [94], the segmentation results of UnmaskFormer, and ground truth. Comparing the performance of UnmaskFormer with the prior state-of-the-art model, Trans4PASS [94], it can be seen that UnmaskFormer obtains notable improvements in handling adverse weather conditions. For instance, in cloudy weather, UnmaskFormer demonstrates an enhanced capability to predict and generate more complete rail track categories. Moreover, in foggy and rainy weather, UnmaskFormer can identify roads more completely compared to the baseline model. Thanks to the deformable designs, UnmaskFormer can yield better panoramic semantic segmentation results. These observations demonstrate the efficacy of UnmaskFormer in challenging environmental scenarios.
| Building | Fence | Other | Pedestrian | Pole | Road Line | Road | Sidewalk | Vegetation | Vehicles | Wall |
| Tr. Sign | Sky | Ground | Bridge | Rail Track | Ground Rail | Tr. Light | Static | Dynamic | Water | Terrain |
E Discussion
Future work. The advancements made in Occlusion-Aware Seamless Segmentation (OASS) through our UnmaskFormer open new avenues for future research. In the future, we envision several potential further directions: 1) User-Interactive Image Editing: We propose exploring user interaction in image editing. Future iterations can integrate user-selected areas to separate amodal layers and employ diffusion processes for completion. These interactive features enhance user-friendliness and expand practical applications. 2) Amodal Optical Flow Fusion: Investigate combining our segmentation approach with amodal optical flow techniques [52] to improve temporal consistency in amodal segmentation. This is useful for dealing with temporal occlusions. Also, it is essential for continuous tracking or monitoring applications, potentially revolutionizing dynamic scene processing.
Limitations and potential solutions. While our UnmaskFormer for OASS showcases significant advancements, we acknowledge certain limitations in our current work, alongside potential approaches to address these challenges: 1) Complex Environment Perception: The current model may not handle the complexity of in-the-wild scenes. Continuous improvement is vital for panoramic image processing and scene understanding. We could enhance robustness through diverse data, integration with other sensors like LiDAR or event cameras, and depth-aware transferring learning [8] to enhance occlusion reasoning. 2) Amodal Data Annotation Challenge: Amodal data annotation for panoramic images is challenging. While our dataset covers diverse images captured in cities located on all continents, we could further explore automated amodal annotation for panoramas, reserving manual annotation for data purification and cleaning purposes. This approach can help scale up the availability of annotated data for training and validation, addressing the scarcity of labeled data in this domain.
Societal impacts. In this study, we have introduced a novel task called Occlusion-Aware Seamless Segmentation (OASS) and established a comprehensive benchmark incorporating various well-known baseline models. We found that these baseline models exhibit limited performance in the OASS task, primarily due to the intricate nature of occlusion-aware segmentation challenges. To address this, we developed UnmaskFormer, a solution that significantly enhances performance on the OASS benchmark, outperforming existing domain adaptation panoramic and panoptic segmentation methods and achieving promising state-of-the-art results. Nevertheless, given the criticality of dependability in deep learning systems for Advanced Driver-Assistance Systems (ADAS), it is important to note that UnmaskFormer may still encounter misclassifications in challenging occluded regions and biased content, potentially leading to erroneous predictions with adverse societal implications.